# Disentangling, Amplifying, and Debiasing: Learning Disentangled Representations for Fair Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) have become essential tools for graph representation learning in various domains, such as social media and healthcare. However, they often suffer from fairness issues due to inherent biases in node attributes and graph structure, leading to unfair predictions. To address these challenges, we propose a novel GNN framework, DAB-GNN, that Disentangles, Amplifies, and deBiases attribute, structure, and potential biases in the GNN mechanism. DAB-GNN employs a disentanglement and amplification module that isolates and amplifies each type of bias through specialized disentanglers, followed by a debiasing module that minimizes the distance between subgroup distributions to ensure fairness. Extensive experiments on five datasets demonstrate that DAB-GNN significantly outperforms ten state-of-the-art competitors in terms of achieving an optimal balance between accuracy and fairness.
@@ 1 Introduction
Background. In the real world, data from various domains such as social media, healthcare, and finance can be represented as graph data (Rozemberczki et al. 2022; Tang and Liu 2009; Yoo et al. 2023). In these graphs, entities (e.g., users) are depicted as nodes, and pairwise relationships between these entities (e.g., friendship) are depicted as edges. This structural representation enables the effective analysis of complex relationships within the graph data.
Recent advancements in graph representation learning have utilized Graph Neural Networks (GNNs) to map nodes into low-dimensional embedding space (Kipf and Welling 2017; Kim, Lee, and Kim 2023; Sharma et al. 2024). Using a message-passing framework, GNNs iteratively aggregate information from a node and its neighbors to produce the final embedding of the node that effectively captures both node attributes and the graph’s structure (Hamilton, Ying, and Leskovec 2017; Kipf and Welling 2017). Consequently, GNNs have demonstrated superior performance in a variety of tasks like node classification and link prediction (Sun et al. 2024; Chamberlain et al. 2023).
Motivation. However, predictions based on node embeddings learned by GNNs can be unfair due to the biases inherent in the input graph data, such as node attributes and graph structure, as well as the biases introduced by the message-passing mechanism of GNNs (Dai and Wang 2021; Wang et al. 2022; Ling et al. 2023; Li et al. 2024; Jiang et al. 2024). Specifically, node attributes may exhibit different distributions across subgroups (e.g., male and female) defined by sensitive attributes (e.g., gender), referred to as attribute bias (Dong et al. 2022). For instance, in job application data, the average income of employees may differ across genders. Additionally, the graph structure itself can be biased, as nodes with similar sensitive attributes tend to form connections, known as structure bias (Dong et al. 2022). For example, on social network platforms, users predominantly form friendships within same sensitive attributes (Rahman et al. 2019; Dai and Wang 2021).
@@
  • Novel Framework: We introduce DAB-GNN, a novel versatile framework that learns disentangled embeddings for fair graph neural networks.
@@ 2 Preliminaries
Fairness in graph mining, particularly in the context of GNNs, is a crucial area of research. Commonly studied notions in FGNN include group fairness and individual fairness (Du et al. 2021). [^2] Another notion, counterfactual fairness, requires that the prediction outcome for a node remains unchanged even if its sensitive attribute was different. This notion often overlaps with the objective of either group fairness or individual fairness.
  • Group Fairness ensures that an algorithm does not produce biased outcomes against minority groups defined by sensitive attributes, such as race and gender.
  • Individual Fairness ensures that an algorithm gives similar outcomes to similar nodes.
In this paper, we focus on group fairness, the most widely studied concept in FGNN research, and review recent studies proposed to ensure group fairness in GNNs.
@@
Detailed equations for calculating each of these metrics can be found in Section A of the supplementary material.
@@
Table 2 presents the results of the fairness metrics. Higher values for attribute bias and structure bias, along with a homophily ratio closer to 1 and neighborhood fairness closer to 0, indicate a higher degree of inherent bias in each dataset. The analysis reveals that biases vary across datasets, with each exhibiting different dominant biases. For example, the NBA dataset shows significant attribute and structure biases, high homophily, and moderate neighborhood fairness, while the Pokec datasets show low attribute and structure biases but high homophily and low neighborhood fairness. These findings support our claim in Section 1, emphasizing the importance of addressing each bias according to its unique characteristics, as biases manifest differently and do not always follow consistent patterns in different datasets.
@@ 3 The Proposed Framework: DAB-GNN
  • Attribute Bias Disentangler (AbDisen): This component leverages a specialized GNN to effectively capture the attribute bias. The process begins with the attribute matrix $\mathbf{X}_{\text{attr}}\in\mathbb{R}^{n\times d}$, which represents the node attributes of the input graph. Using $\mathbf{X}_{\text{attr}}$, we construct a $k$-Nearest Neighbors ($k$-NN) graph based on the Euclidean distance between node attributes. [^3] For sensitivity analysis of the performance of DAB-GNN across different $k$ values, see Section C of the supplementary material. The distance between two nodes $i$ and $j$ is quantified as $\texttt{D}_{e}(i,j)=\sqrt{\sum_{r=1}^{d}(x_{ir}-x_{jr})^{2}}$, where ${x}_{i}$ and ${x}_{j}$ represent the attribute vectors for nodes $i$ and $j$ in $\mathbf{X}_{\text{attr}}$, respectively. Based on $\texttt{D}_{e}(i,\cdot)$, the $k$ nearest neighbors for each node $i$ are identified. The $k$-NN graph is then represented by the adjacency matrix $\mathbf{A}_{\text{attr}}\in\mathbb{R}^{n\times n}$:
@@
    $\mathbf{H}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}attr}}}^{(l+1)}=\sigma\left(\mathbf{A}_{\text{attr}}\cdot\mathbf{H}_{\text{attr}}^{(l)}\cdot\mathbf{W}_{\text{attr}}^{(l)}+\mathbf{b}\right),$ (2)
@@
    $\mathbf{H}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}stru}}}^{(l+1)}=\sigma\left(\mathbf{A}_{\text{stru}}\cdot\mathbf{H}_{\text{stru}}^{(l)}\cdot\mathbf{W}_{\text{stru}}^{(l)}+\mathbf{b}\right),$ (3)
@@
    $\mathbf{H}_{\text{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}ent}}}=\texttt{MLP}\left([\mathbf{H}_{\text{attr}}\,|\,\mathbf{H}_{\text{stru}}]\right),$ (4)
@@
    $\mathbf{H}_{\text{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}pot}}}=\mathbf{H}_{\text{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}ent}}}-\mathbf{H}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}attr}}}-\mathbf{H}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}stru}}}.$ (5)
@@
Discussions
Our primary contribution lies not in creating a new GNN architecture, but in developing a versatile framework for achieving fairness in GNNs, which can easily integrate with any existing GNN architecture. The DAB-GNN framework introduces a unique approach for disentangling and independently addressing different types of biases, offering a fresh perspective on ensuring fairness in GNNs. As demonstrated in the following section, even a simple implementation of DAB-GNN outperforms state-of-the-art FGNN methods, effectively balancing accuracy and fairness. This underscores the broad applicability of DAB-GNN across various domains where fairness in graph-based models is crucial.
Table 4: Accuracy and fairness results of DAB-GNN and competitors across five real-world datasets. ($\uparrow$) and ($\downarrow$) mean higher and lower values are better, respectively; ‘o.o.m’ denotes ‘out of memory’ and values in parentheses represent the standard deviation.
@@
NBA ACC ($\uparrow$) 57.97 (2.45) 58.73 (1.29) 60.76 (5.25) 63.29 (5.88) 69.11 (2.61) 65.57 (6.86) 60.51 (6.86) 57.98 (1.24) 60.76 (5.13) 70.63 (2.18) 58.73 (1.29) 66.58 (4.43) 71.39 (2.35)
AUC ($\uparrow$) 63.75 (1.53) 63.33 (0.80) 74.91 (2.29) 70.75 (8.67) 71.82 (1.69) 79.96 (2.16) 67.06 (10.51) 60.04 (1.11) 79.33 (0.96) 73.26 (0.99) 63.33 (0.80) 71.72 (2.51) 80.56 (1.18)
F1 ($\uparrow$) 61.55 (2.26) 62.00 (1.21) 70.69 (1.83) 66.86 (12.26) 74.99 (2.19) 72.93 (2.40) 68.81 (4.04) 65.08 (1.88) 70.50 (0.47) 74.06 (2.38) 62.00 (1.21) 74.21 (2.10) 73.51 (3.22)
SP ($\downarrow$) 32.94 (5.91) 32.83 (0.97) 6.39 (5.35) 9.82 (8.14) 8.98 (3.60) 7.82 (7.14) 0.00 (0.00) 20.53 (3.30) 4.55 (4.45) 4.03 (1.88) 32.83 (0.97) 12.96 (4.86) 1.12 (0.73)
EO ($\downarrow$) 33.68 (6.45) 35.95 (4.41) 10.14 (9.36) 8.60 (10.16) 4.39 (2.72) 13.28 (13.60) 0.00 (0.00) 21.94 (6.20) 1.77 (1.75) 13.56 (5.82) 35.95 (4.41) 2.34 (1.81) 0.80 (0.28)
Recidivism ACC ($\uparrow$) 84.18 (0.75) 83.73 (0.46) 84.50 (0.49) 79.94 (0.41) 78.18 (7.16) 83.64 (0.91) 86.79 (0.41) 78.32 (1.77) 84.49 (1.11) 85.41 (0.33) 81.28 (0.47) 86.59 (0.47) 89.99 (1.50)
AUC ($\uparrow$) 86.90 (0.23) 86.84 (0.18) 89.05 (0.48) 81.23 (0.21) 83.62 (5.68) 84.38 (1.24) 87.07 (0.46) 81.30 (0.48) 89.13 (0.35) 89.48 (0.31) 83.23 (0.20) 89.08 (0.63) 93.41 (0.83)
F1 ($\uparrow$) 78.65 (0.84) 78.10 (0.53) 79.77 (0.47) 69.77 (0.45) 73.16 (6.85) 76.89 (1.44) 80.63 (0.43) 71.18 (1.48) 79.82 (0.95) 79.48 (0.56) 75.91 (0.51) 80.87 (0.93) 86.31 (1.73)
SP ($\downarrow$) 7.79 (0.54) 8.13 (0.25) 6.64 (0.59) 3.69 (0.43) 10.89 (2.13) 5.42 (0.72) 5.73 (0.19) 5.81 (1.06) 9.24 (0.42) 6.13 (0.56) 1.43 (0.26) 5.65 (0.82) 0.73 (0.49)
EO ($\downarrow$) 5.23 (0.46) 5.65 (0.25) 3.16 (0.88) 2.97 (0.41) 7.62 (2.99) 3.92 (0.87) 3.41 (0.24) 4.11 (0.74) 4.61 (0.31) 4.14 (0.40) 2.92 (0.25) 3.59 (1.54) 0.90 (0.65)
Credit ACC ($\uparrow$) 73.57 (0.07) 73.92 (0.04) 73.99 (3.76) 73.43 (0.05) 74.77 (0.81) 77.92 (0.07) 76.00 (2.16) o.o.m 74.60 (1.04) 63.96 (0.20) 73.21 (1.06) 77.60 (0.50) 78.19 (0.68)
AUC ($\uparrow$) 73.48 (0.62) 73.40 (0.14) 64.19 (10.31) 72.14 (0.21) 72.30 (0.51) 68.67 (1.39) 65.72 (2.04) o.o.m 71.91 (0.23) 66.90 (0.04) 70.10 (3.99) 71.57 (0.15) 71.41 (1.96)
F1 ($\uparrow$) 81.87 (0.06) 82.16 (0.04) 83.08 (4.15) 81.70 (0.04) 82.99 (0.78) 87.48 (0.11) 85.15 (1.94) o.o.m 82.76 (0.94) 73.95 (0.22) 82.03 (0.16) 87.23 (0.53) 87.39 (0.49)
SP ($\downarrow$) 13.88 (1.80) 12.18 (0.19) 3.17 (2.99) 11.60 (0.09) 7.98 (5.01) 0.40 (0.56) 11.70 (4.76) o.o.m 11.76 (0.93) 19.19 (0.80) 1.39 (2.28) 0.69 (0.15) 0.44 (0.47)
EO ($\downarrow$) 11.68 (1.80) 10.04 (0.15) 1.73 (2.02) 9.30 (0.13) 6.09 (4.78) 0.16 (0.19) 8.51 (4.63) o.o.m 9.15 (1.04) 22.66 (0.87) 1.83 (2.06) 0.66 (0.25) 0.45 (0.57)
Pokec_n ACC ($\uparrow$) 66.97 (0.30) 65.27 (2.71) 63.56 (4.34) 67.86 (1.18) o.o.m 69.51 (0.31) o.o.m o.o.m 55.69 (0.39) o.o.m 66.54 (0.61) 65.69 (1.19) 67.18 (1.16)
AUC ($\uparrow$) 72.73 (0.41) 70.74 (1.67) 67.10 (4.52) 73.92 (0.16) o.o.m 73.99 (0.34) o.o.m o.o.m 58.99 (0.27) o.o.m 71.76 (0.55) 72.89 (0.60) 73.68 (1.32)
F1 ($\uparrow$) 65.70 (0.36) 64.91 (1.29) 59.79 (5.46) 66.25 (1.31) o.o.m 66.01 (0.70) o.o.m o.o.m 52.36 (0.33) o.o.m 65.91 (0.55) 67.44 (0.75) 62.34 (1.83)
SP ($\downarrow$) 7.90 (3.95) 17.19 (3.77) 3.28 (0.69) 1.20 (0.73) o.o.m 2.77 (1.37) o.o.m o.o.m 6.78 (0.86) o.o.m 14.97 (2.19) 2.40 (1.88) 0.71 (0.37)
EO ($\downarrow$) 7.09 (3.95) 14.88 (2.01) 5.05 (1.00) 1.23 (0.53) o.o.m 3.38 (2.23) o.o.m o.o.m 5.96 (0.95) o.o.m 11.38 (0.39) 1.64 (1.55) 1.09 (0.93)
Pokec_z ACC ($\uparrow$) 64.92 (0.13) 65.40 (0.20) 62.97 (4.69) 65.71 (0.48) o.o.m 63.38 (5.77) o.o.m o.o.m 58.38 (0.41) o.o.m 64.39 (0.35) 62.21 (0.96) 68.56 (0.79)
AUC ($\uparrow$) 70.03 (0.22) 69.84 (0.15) 65.81 (5.94) 70.57 (0.24) o.o.m 68.99 (3.48) o.o.m o.o.m 61.20 (0.50) o.o.m 69.08 (0.31) 68.81 (0.52) 74.85 (0.78)
F1 ($\uparrow$) 65.48 (0.22) 65.08 (0.83) 64.47 (2.65) 65.00 (1.08) o.o.m 67.31 (1.39) o.o.m o.o.m 58.13 (0.56) o.o.m 65.45 (0.40) 65.37 (1.69) 67.94 (1.81)
SP ($\downarrow$) 7.27 (1.09) 10.91 (4.38) 4.79 (4.56) 5.03 (0.47) o.o.m 5.04 (3.87) o.o.m o.o.m 6.13 (1.00) o.o.m 12.18 (1.01) 0.96 (0.39) 0.67 (0.64)
EO ($\downarrow$) 4.05 (1.26) 7.88 (2.14) 3.65 (3.23) 1.24 (0.17) o.o.m 3.06 (2.88) o.o.m o.o.m 4.96 (1.31) o.o.m 7.14 (0.86) 1.64 (1.35) 0.73 (0.35)
Table 5: The ablation studies for DAB-GNN
Metrics (a) Disen. and Amp. Module (b) Debiasing Module DAB-GNN
NBA ACC ($\uparrow$) 70.38 (2.35) 70.89 (1.13) 71.65 (1.01) 69.11 (3.26) 69.37 (1.86) 71.39 (2.35)
AUC ($\uparrow$) 79.51 (1.65) 81.13 (1.41) 81.86 (0.47) 77.68 (4.38) 81.14 (2.72) 80.56 (1.18)
F1 ($\uparrow$) 73.14 (2.93) 74.07 (1.66) 73.36 (1.96) 70.48 (3.19) 72.05 (2.66) 73.51 (3.22)
SP ($\downarrow$) 6.08 (5.22) 5.97 (3.20) 6.52 (5.97) 9.50 (11.61) 6.76 (6.74) 1.12 (0.73)
EO ($\downarrow$) 8.66 (7.74) 8.26 (6.47) 12.25 (8.12) 12.65 (12.39) 8.03 (6.47) 0.80 (0.28)
Recidivism ACC ($\uparrow$) 88.82 (0.42) 91.59 (0.35) 91.25 (0.42) 91.63 (0.51) 91.01 (0.71) 89.99 (1.50)
AUC ($\uparrow$) 92.38 (0.44) 94.13 (0.39) 94.07 (0.30) 94.24 (0.26) 94.04 (0.79) 93.41 (0.83)
F1 ($\uparrow$) 84.20 (0.61) 88.17 (0.41) 87.72 (0.56) 87.73 (0.90) 87.53 (1.07) 86.31 (1.73)
SP ($\downarrow$) 1.13 (0.59) 1.21 (0.94) 1.03 (0.31) 3.30 (0.41) 0.86 (0.60) 0.73 (0.49)
EO ($\downarrow$) 0.77 (0.38) 2.18 (1.23) 1.45 (0.92) 2.26 (0.63) 1.61 (0.74) 0.90 (0.65)
Credit ACC ($\uparrow$) 74.22 (3.17) 75.06 (2.52) 75.75 (2.68) 76.47 (2.65) 74.10 (3.75) 78.19 (0.68)
AUC ($\uparrow$) 72.86 (1.50) 68.45 (5.63) 72.05 (2.48) 64.08 (4.38) 69.66 (4.70) 71.41 (1.96)
F1 ($\uparrow$) 82.84 (3.27) 84.04 (3.06) 84.80 (3.13) 85.93 (3.05) 83.16 (3.69) 87.39 (0.49)
SP ($\downarrow$) 6.89 (5.17) 5.68 (6.25) 4.73 (5.19) 2.59 (4.95) 5.98 (5.08) 0.44 (0.47)
EO ($\downarrow$) 5.67 (4.40) 4.90 (5.26) 3.76 (4.23) 2.30 (4.45) 5.12 (4.38) 0.45 (0.57)
Pokec_n ACC ($\uparrow$) 61.18 (6.43) 67.33 (0.57) 66.50 (1.22) 66.79 (1.21) 67.75 (0.71) 67.18 (1.16)
AUC ($\uparrow$) 67.87 (5.35) 72.89 (1.24) 72.96 (1.57) 73.62 (1.76) 72.82 (1.06) 73.68 (1.32)
F1 ($\uparrow$) 58.22 (6.39) 61.48 (2.09) 63.85 (1.58) 64.54 (0.75) 62.78 (1.01) 62.34 (1.83)
SP ($\downarrow$) 2.03 (1.41) 2.21 (2.73) 2.04 (0.91) 4.05 (1.66) 1.38 (0.94) 0.71 (0.37)
EO ($\downarrow$) 2.57 (2.20) 3.30 (1.74) 2.64 (0.96) 5.40 (1.96) 3.61 (1.75) 1.09 (0.93)
Pokec_z ACC ($\uparrow$) 67.32 (2.03) 68.92 (0.55) 68.82 (0.54) 69.02 (0.54) 67.99 (0.94) 68.56 (0.79)
AUC ($\uparrow$) 73.70 (1.72) 75.14 (0.92) 74.49 (1.06) 75.39 (0.38) 75.05 (0.91) 74.85 (0.78)
F1 ($\uparrow$) 65.89 (3.21) 67.60 (1.20) 67.16 (1.56) 68.30 (0.76) 69.17 (1.89) 67.94 (1.81)
SP ($\downarrow$) 3.55 (3.28) 1.88 (1.50) 4.82 (5.00) 4.59 (0.96) 2.43 (1.11) 0.67 (0.64)
EO ($\downarrow$) 3.70 (2.65) 1.94 (1.40) 3.97 (3.97) 3.61 (0.57) 2.51 (1.30) 0.73 (0.35)

### v2

@@ Abstract
Graph Neural Networks (GNNs) have become essential tools for graph representation learning in various domains, such as social media and healthcare. However, they often suffer from fairness issues due to inherent biases in node attributes and graph structure, leading to unfair predictions. To address these challenges, we propose a novel GNN framework, DAB-GNN, that Disentangles, Amplifies, and deBiases attribute, structure, and potential biases in the GNN mechanism. DAB-GNN employs a disentanglement and amplification module that isolates and amplifies each type of bias through specialized disentanglers, followed by a debiasing module that minimizes the distance between subgroup distributions. Extensive experiments on five datasets demonstrate that DAB-GNN significantly outperforms ten state-of-the-art competitors in terms of achieving an optimal balance between accuracy and fairness.
Codehttps://github.com/Bigdasgit/DAB-GNN
@@ 1 Introduction
Background. In the real world, data from various domains such as social media, healthcare, and finance can be represented as graph data (Rozemberczki et al. 2022; Tang and Liu 2009; Yoo et al. 2023; Kim et al. 2024). In these graphs, entities (e.g., users) are depicted as nodes, and pairwise relationships between these entities (e.g., friendship) are depicted as edges. This structural representation enables the effective analysis of complex relationships within the graph data.
Recent advancements in graph representation learning have utilized Graph Neural Networks (GNNs) to map nodes into low-dimensional embedding space (Kipf and Welling 2017; Kim, Lee, and Kim 2023; Sharma et al. 2024). Using a message-passing framework, GNNs iteratively aggregate information from a node and its neighbors to produce the final embedding of the node that effectively captures both node attributes and the graph’s structure (Hamilton, Ying, and Leskovec 2017; Kipf and Welling 2017). Consequently, GNNs have demonstrated superior performance in a variety of tasks like node classification and link prediction (Sun et al. 2024; Chamberlain et al. 2023; Kim, Lee, and Kim 2024).
Motivation. However, predictions based on node embeddings learned by GNNs can be unfair due to the biases inherent in the input graph data, such as node attributes and graph structure, as well as the biases introduced by the message-passing mechanism of GNNs (Dai and Wang 2021; Wang et al. 2022; Ling et al. 2023; Li et al. 2024; Jiang et al. 2024; Neo et al. 2024). Specifically, node attributes may exhibit different distributions across subgroups (e.g., male and female) defined by sensitive attributes (e.g., gender), referred to as attribute bias (Dong et al. 2022). For instance, in job application data, the average income of employees may differ across genders. Additionally, the graph structure itself can be biased, as nodes with similar sensitive attributes tend to form connections, known as structure bias (Dong et al. 2022). For example, on social network platforms, users predominantly form friendships within same sensitive attributes (Rahman et al. 2019; Dai and Wang 2021).
@@
  • Novel Framework: We introduce DAB-GNN, a framework that learns disentangled embeddings for fair GNNs.
@@ 2 Preliminaries
Fairness in graph mining, particularly in the context of GNNs, is a crucial area of research. Commonly studied notions in FGNN include group fairness and individual fairness (Du et al. 2021). Group fairness ensures that an algorithm does not produce biased outcomes against minority groups defined by sensitive attributes On the other hand, individual fairness ensures that an algorithm gives similar outcomes to similar nodes. We focus on group fairness, the most widely studied concept in FGNN research, and review recent studies proposed to ensure group fairness in GNNs.
@@
Detailed equations for calculating each of these metrics can be found in the online appendix at https://github.com/Bigdasgit/DAB-GNN.
@@
Table 2 presents the results of the fairness metrics. Higher values for attribute bias and structure bias, along with a homophily ratio closer to 1 and neighborhood fairness closer to 0, indicate a higher degree of inherent bias in each dataset. The analysis reveals that biases vary across datasets, with each exhibiting different dominant biases. For example, the NBA dataset shows significant attribute and structure biases, high homophily, and moderate neighborhood fairness, while the Pokec datasets show low attribute and structure biases but high homophily and low neighborhood fairness These findings support our claim in Section 1, emphasizing the importance of addressing each bias according to its unique characteristics, as biases manifest differently and do not always follow consistent patterns in different datasets.
@@ 3 The Proposed Framework: DAB-GNN
Table 3: Notations used in this paper.
  • Attribute Bias Disentangler (AbDisen): This component leverages a specialized GNN to effectively capture the attribute bias. The process begins with the attribute matrix $\mathbf{X}_{\text{attr}}\in\mathbb{R}^{n\times d}$, which represents the node attributes of the input graph. Using $\mathbf{X}_{\text{attr}}$, we construct a $k$-Nearest Neighbors ($k$-NN) graph based on the Euclidean distance between node attributes. The distance between two nodes $i$ and $j$ is quantified as $\texttt{D}_{e}(i,j)=\sqrt{\sum_{r=1}^{d}(x_{ir}-x_{jr})^{2}}$, where ${x}_{i}$ and ${x}_{j}$ represent the attribute vectors for nodes $i$ and $j$ in $\mathbf{X}_{\text{attr}}$, respectively. Based on $\texttt{D}_{e}(i,\cdot)$, the $k$ nearest neighbors for each node $i$ are identified. The $k$-NN graph is then represented by the adjacency matrix $\mathbf{A}_{\text{attr}}\in\mathbb{R}^{n\times n}$:
@@
    $\mathbf{H}_{\text{attr}}^{(l+1)}=\sigma\left(\mathbf{A}_{\text{attr}}\cdot\mathbf{H}_{\text{attr}}^{(l)}\cdot\mathbf{W}_{\text{attr}}^{(l)}+\mathbf{b}\right),$ (2)
@@
    $\mathbf{H}_{\text{stru}}^{(l+1)}=\sigma\left(\mathbf{A}_{\text{stru}}\cdot\mathbf{H}_{\text{stru}}^{(l)}\cdot\mathbf{W}_{\text{stru}}^{(l)}+\mathbf{b}\right),$ (3)
@@
    $\mathbf{H}_{\text{ent}}=\texttt{MLP}\left([\mathbf{H}_{\text{attr}}\,|\,\mathbf{H}_{\text{stru}}]\right),$ (4)
@@
    $\mathbf{H}_{\text{pot}}=\mathbf{H}_{\text{ent}}-\mathbf{H}_{\text{attr}}-\mathbf{H}_{\text{stru}}.$ (5)
@@
NBA ACC ($\uparrow$) 57.97 58.73 60.76 63.29 69.11 65.57 60.51 57.98 60.76 70.63 58.73 66.58 71.39
AUC ($\uparrow$) 63.75 63.33 74.91 70.75 71.82 79.96 67.06 60.04 79.33 73.26 63.33 71.72 80.56
F1 ($\uparrow$) 61.55 62.00 70.69 66.86 74.99 72.93 68.81 65.08 70.50 74.06 62.00 74.21 73.51
SP ($\downarrow$) 32.94 32.83 6.39 9.82 8.98 7.82 0.00 20.53 4.55 4.03 32.83 12.96 1.12
EO ($\downarrow$) 33.68 35.95 10.14 8.60 4.39 13.28 0.00 21.94 1.77 13.56 35.95 2.34 0.80
Recidivism ACC ($\uparrow$) 84.18 83.73 84.50 79.94 78.18 83.64 86.79 78.32 84.49 85.41 81.28 86.59 89.99
AUC ($\uparrow$) 86.90 86.84 89.05 81.23 83.62 84.38 87.07 81.30 89.13 89.48 83.23 89.08 93.41
F1 ($\uparrow$) 78.65 78.10 79.77 69.77 73.16 76.89 80.63 71.18 79.82 79.48 75.91 80.87 86.31
SP ($\downarrow$) 7.79 8.13 6.64 3.69 10.89 5.42 5.73 5.81 9.24 6.13 1.43 5.65 0.73
EO ($\downarrow$) 5.23 5.65 3.16 2.97 7.62 3.92 3.41 4.11 4.61 4.14 2.92 3.59 0.90
Credit ACC ($\uparrow$) 73.57 73.92 73.99 73.43 74.77 77.92 76.00 o.o.m 74.60 63.96 73.21 77.60 78.19
AUC ($\uparrow$) 73.48 73.40 64.19 72.14 72.30 68.67 65.72 o.o.m 71.91 66.90 70.10 71.57 71.41
F1 ($\uparrow$) 81.87 82.16 83.08 81.70 82.99 87.48 85.15 o.o.m 82.76 73.95 82.03 87.23 87.39
SP ($\downarrow$) 13.88 12.18 3.17 11.60 7.98 0.40 11.70 o.o.m 11.76 19.19 1.39 0.69 0.44
EO ($\downarrow$) 11.68 10.04 1.73 9.30 6.09 0.16 8.51 o.o.m 9.15 22.66 1.83 0.66 0.45
Pokec_n ACC ($\uparrow$) 66.97 65.27 63.56 67.86 o.o.m 69.51 o.o.m o.o.m 55.69 o.o.m 66.54 65.69 67.18
AUC ($\uparrow$) 72.73 70.74 67.10 73.92 o.o.m 73.99 o.o.m o.o.m 58.99 o.o.m 71.76 72.89 73.68
F1 ($\uparrow$) 65.70 64.91 59.79 66.25 o.o.m 66.01 o.o.m o.o.m 52.36 o.o.m 65.91 67.44 62.34
SP ($\downarrow$) 7.90 17.19 3.28 1.20 o.o.m 2.77 o.o.m o.o.m 6.78 o.o.m 14.97 2.40 0.71
EO ($\downarrow$) 7.09 14.88 5.05 1.23 o.o.m 3.38 o.o.m o.o.m 5.96 o.o.m 11.38 1.64 1.09
Pokec_z ACC ($\uparrow$) 64.92 65.40 62.97 65.71 o.o.m 63.38 o.o.m o.o.m 58.38 o.o.m 64.39 62.21 68.56
AUC ($\uparrow$) 70.03 69.84 65.81 70.57 o.o.m 68.99 o.o.m o.o.m 61.20 o.o.m 69.08 68.81 74.85
F1 ($\uparrow$) 65.48 65.08 64.47 65.00 o.o.m 67.31 o.o.m o.o.m 58.13 o.o.m 65.45 65.37 67.94
SP ($\downarrow$) 7.27 10.91 4.79 5.03 o.o.m 5.04 o.o.m o.o.m 6.13 o.o.m 12.18 0.96 0.67
EO ($\downarrow$) 4.05 7.88 3.65 1.24 o.o.m 3.06 o.o.m o.o.m 4.96 o.o.m 7.14 1.64 0.73
Table 4: Accuracy and fairness results of DAB-GNN and competitors across five real-world datasets. ($\uparrow$) and ($\downarrow$) mean higher and lower values are better, respectively; ‘o.o.m’ denotes ‘out of memory.’